OpenStudy (anonymous):

-2y=6(5-3x) how do i turn this into slope intercept

OpenStudy (anonymous):

plz reply

OpenStudy (anonymous):

Release parathentles: -2y=6*5-6*3x=30-18x -2y=30-18x y=30/(-2)-18/(-2)x=-15+9x y=9x-15